# The Jordan canonical form of the Fréchet derivative of a matrix function

### v1

@@ 4.2 Equal eigenvalues
Proof. We only prove the nontrivial case kd<m+n1kd<m+n-1. Note by Lemma 4.7 that $d$ is equal to the local degree of $p(x,y)$ at $(\lambda,\lambda)$. We may assume without loss of generality that that $p(\lambda,\lambda)=0=\hat{\partial}_{w}f(\lambda)$. Observe that, for some $f_{\ell},g_{\ell}\in\mathbb{F}$,
$p(x,y)=\sum_{\ell=d}^{D}f_{\ell}H_{\ell}(x,y)\Rightarrow p(x,y)^{k}=(f_{d})^{k}H_{d}(x,y)^{k}+\sum_{\ell=kd+1}^{kD}g_{\ell}H_{\ell}(x,y).$
Thus, if $p(x,y)$ has local degree $d$ then $p(x,y)^{k}$ has local degree $kd$. Set now for simplicity $Z_{m-1}:=P(J_{m}(\lambda),J_{n}(\lambda))-p(\lambda,\lambda)I_{mn}$. By Theorem 4.9 and the observation above, $Z_{m-1}^{k}$ is similar to $H_{kd}(N_{m},N_{n})$. Hence, by repeated applications of Proposition 4.11, we have $\dim\ker(Z_{m-1}^{k})=\mathfrak{N}_{k}.$
To conclude the proof, we invoke the classical result [14] that the number of Jordan blocks of size exactly $s$ is determined by the second differences of the nullity sequence $(\mathfrak{N}_{k})_{k}$.
@@
Let us revisit Example 3.3 to verify that Theorem 4.12 makes the same predictions as Theorem 3.1. Clearly $d=1$ and $m=n=2$. Let us now use Proposition 4.11. For $k=1$,
$\mathfrak{R}(1)=\rho(1,2,0,2)+\rho(2,1,0,2)=(2-1)+(1-0)=2.$
For $k=2$,
$\mathfrak{R}(2)=\rho(1,1,1,2)=1-0=1.$
For the other values of $k$ $\mathfrak{R}(kd)$ is trivial, and the sequence $(\mathfrak{N}_{k})_{k\in\mathbb{N}}$ in Theorem 4.12 is
@@
We can then apply Theorem 4.12 with $m=2,n=3$. The only notrivial value of $\mathfrak{R}$ that we need to compute is
$\mathfrak{R}(2)=\rho(1,2,0,2)+\rho(2,1,1,2)=(2-1)+(1-0)=2$
@@
In many examples (especially if $d$ is small compared to $m,n$), the lower bound $\mathfrak{b}$ of Proposition 3.11 is attained by Fréchet derivatives; since the expression for $\mathfrak{b}$ is significantly simpler than that for $\mathfrak{R}$ in Proposition 4.11, one might be tempted to get a simpler formula by conjecturing that the lower bound $\mathfrak{b}$ is always attained by the polynomials $H_{k}$. However, this is false in general, as Example 4.15 below demonstrates.
@@
Let $p(x,y)=H_{4}(x,y)$, corresponding to the formal Fréchet derivative of $f(w)=w^{5}$. Take as matrix arguments $X=Y=J_{4}(0))$; the local degree at $(0,0)$ is $d=4$. We have $\mathfrak{R}(4k)=0$ for $k\geq 2$, and
$\mathfrak{R}(4)=\rho(1,3,1,4)+\rho(2,2,2,4)+\rho(3,1,3,4)=1+1+1=3,$
yielding
@@
We emphasize that for this example the lower bound of Proposition 3.11 is not attained:
@@ 5 Conclusions and future outlook
We studied two nontrivial problems of combinatorial nature, the Fréchet-Jordan problem and the more general bivariate Jordan problem. We fully solved the former. For the latter, we gave some partial results, complementing the existing literature. A full solution of the bivariate Jordan problem remains an open problem, and indeed not an easy one in the author’s opinion. Extending our analysis to fields of prime characteristic is also a possible direction for future research.
@@ 6 Acknowledgements
The late Nick Higham suggested the Fréchet-Jordan problem to me when I was a postdoc in his group. Nick himself, Bruno Iannazzo, and Samuel Relton (as well as Giovanni Barbarino, a few years later) shared comments on an earlier version of this document, that did not attempt to fully solve the problem and did not go much beyond Theorem 3.1. This earlier version was in fact submitted, and an anonymous reviewer pointed me to [9], which in turn cites [11]. My discovery of [11] implied that some of the results that I believed new were not, and I decided to not revise that version. I retracted it, planning to only resubmit after having solved the Fréchet-Jordan problem for good.

### v2

@@ 4.2 Equal eigenvalues
Proof. The statement follows by Theorem 4.9, the discussion above, and [15].
@@
Let us revisit Example 3.3 to verify that Theorem 4.12 makes the same predictions as Theorem 3.1. Clearly $d=1$ and $m=n=2$. Let us now use Theorem 4.12. For $s\geq 3$, $\mathfrak{N}_{s}=4$ is trivial. For $s=1$, we must sum the ranks of the Toeplitz matrices
$R_{2}=\begin{bmatrix}1&1\end{bmatrix},\qquad R_{3}=\begin{bmatrix}1\\ 1\end{bmatrix}$
to obtain $\mathfrak{N}_{1}=4-1-1=2.$ Similarly, when $s=2$ we subtract from $mn=4$ the rank of $R_{3}=2$, and hence $\mathfrak{N}_{2}=3$.
We conclude that the sequence $(\mathfrak{N}_{s})_{s\in\mathbb{N}}$ in Theorem 4.12 is
@@
We can then apply Theorem 4.12 with $m=2,n=3,d=2$. The only value $s$ for which the computation of $\mathfrak{N}_{s}$ is not trivial is $s=1$. In this case, we subtract from $mn=6$ the ranks of
$R_{3}=\begin{bmatrix}1&1\end{bmatrix},\qquad R_{4}=\begin{bmatrix}1\\ 1\end{bmatrix}$.
@@
Let $p(x,y)=H_{4}(x,y)$, corresponding to the formal Fréchet derivative of $f(w)=w^{5}$. Take as matrix arguments $X=Y=J_{4}(0))$; the local degree at $(0,0)$ is $d=4$. We have $\mathfrak{N}_{s}=16$ for $s\geq 2$. When $s=1$, the quantity $mn-\mathfrak{N}_{1}$ equals to the sum of the ranks of
$R_{5}=\begin{bmatrix}1&1&1\end{bmatrix},\qquad R_{6}=\begin{bmatrix}1&1\\ 1&1\end{bmatrix},\qquad R_{7}=\begin{bmatrix}1\\ 1\\ 1\end{bmatrix}$.
This yields
@@
Due to the nonmaximal rank of $R_{6}$, the lower bound of Proposition 3.11 is not attained:
@@
4.3 Ranks of $R_{k}$: A discussion and an open problem
Theorem 4.12 gives an implicit solution to the Fréchet-Jordan problem, but it is in some sense less satisfactory than Theorem 4.4 in that it does not provide a closed formula in terms of $n,m,d$. Instead, it reduces the problem to that of computing the ranks of certain integer Toeplitz matrices. Several remarks from different angles are appropriate.
  1. 1.
    Computational viewpoint. Fix $m,n,d,\ell$. The matrix $H_{d}(N_{m},N_{m})^{\ell}$ has size $mn$. Instead, for $m\leq n$, there are at most $O(n)$ matrices $R_{k}$ having size at most $O(m)$. Moreover, by item 5 in Proposition 4.11, it suffices to compute the ranks of about half of them. Thus, computing and summing the ranks of $R_{k}$ is computationally more appealing than directly computing the rank of $H_{d}^{\ell}$.
  2. 2.
    Linear algebraic viewpoint. If (having fixed $m,n,\ell,d$) every matrix $R_{k}$ has maximal rank $\min(u_{k},u_{k-\ell d})$, then the nullity of the matrix $H_{d}^{\ell}(N_{m},N_{n})$ achieves the lower bound $\mathfrak{b}$ in Proposition 3.11 (having replaced $d$ with $\ell d$ in the definition of $\delta$, and thus of $\mathfrak{b}$, in the statement of Proposition 3.11). Indeed, in that case, $\mathrm{rank}(H_{d}(N_{m},N_{n})^{\ell})$ is equal to
    $\sum_{k=\ell d+1}^{m+n-1}\min(u_{k},u_{k-\ell d})=mn-\sum_{k=\ell d+1}^{m+n-1}\max(0,u_{k}-u_{k-\ell d})=mn-\mathfrak{b}(m,n,\ell d),$
    where, mutatis mutandis, the last step can be proved by the same argument as in the proof of the lower bound in Proposition 3.11. This is a typical situation, i.e., it is what happens for most quadruples $(m,n,d,\ell)$. Nevertheless, as we already have seen in Example 4.15, a loss of rank in some $R_{k}$ can happen, provided there are nontrivial solutions to the Diophantine linear equations prescribed by $R_{k}$. Below, we provide more insight via Proposition 4.16, Example 4.17, and Example 4.18.
  3. 3.
    Ring theoretical viewpoint. The nullity of $H_{d}(N_{m},N_{n})^{\ell}$ can also be interpreted in terms of bivariate polynomial ideals. Denote by $\mathcal{R}$ the quotient ring $\mathbb{F}[x,y]/\langle x^{m},y^{n}\rangle$. By Remark 1.1, we can interpret $H_{d}(N_{m},N_{n})^{\ell}\in\mathcal{K}_{N_{m},N_{n}}$ as the equivalence class $h_{d}(x,y)^{\ell}]=[h_{d}(x,y)]^{\ell}\in\mathcal{R}$. Therefore, the ring isomorphism in Remark 1.1 induces the vector space isomorphism
    $\ker H_{d}(N_{m},N_{n})^{\ell}\cong\mathcal{I}_{\ell}:=\left\{[p(x,y)]\in\mathcal{R}\ \mid\ p(x,y)h_{d}(x,y)^{\ell}\equiv 0\bmod\langle x^{m},y^{n}\rangle\right\}.$
    The set $\mathcal{I}_{\ell}$ is the annihilator ideal of $h_{d}(x,y)]^{\ell}$. We omit the straightforward verification that it is indeed an ideal of $\mathcal{R}$; note also that, although we have emphasized its dependence on $\ell$, it is clear that $\mathcal{I}_{\ell}$ also depends on $m,n,$ and $d$. If $\ell_{0}$ is the smallest integer such that $d\ell_{0}\geq m+n-1$, we have the ascending chain of ideals
    $\langle[0]\rangle=\mathcal{I}_{0}\subsetneq\mathcal{I}_{1}\subseteq\dots\subseteq\mathcal{I}_{\ell_{0}}=\mathcal{R}=\mathcal{I}_{\ell_{0}+1}=\dots$
    Given $(m,n,d)$, the Jordan form of $H_{d}(N_{m},N_{n})$ is thus completely determined by the sequence $\dim_{\mathbb{F}}\mathcal{I}_{\ell}=\dim_{\mathbb{F}}\ker H_{d}(N_{m},N_{n})^{\ell}$.
We now give, in Proposition 4.16, a sufficient condition for $R_{k}$ to lose rank. By item 5 in Proposition 4.11, we may without loss of generality assume that $R_{k}$ has at least as many rows as columns. It is based on the insight that, for appropriate values of $m,n,d,\ell$, then $(x-y)^{\ell}$ belongs to the annihilator ideal of $h_{d}(x,y)]^{\ell}$.
Proposition 4.16
Fix $m\leq n,d,\ell$ and consider the matrices $R_{k}$ as in Definition 4.10. Suppose with no loss of generality that $\lceil\frac{m+n+\ell d}{2}\rceil\leq k\leq m+n-1$, so that $u_{k-\ell d}\geq u_{k}$, i.e., $R_{k}$ does not have more columns than rows. Denote by $g_{k}:=c_{k}\bmod d+1$ the remainder of $c_{k}$ in the Euclidean division by $d+1$, so that $0\leq g_{k}\leq d$. If (1) $u_{k}>\ell$ and (2) $0<g_{k}<d-u_{k-\ell d}+2$, then $R_{k}$ does not have full rank, i.e., $\mathrm{rank}(R_{k})<u_{k}$.
Proof. Let $0\neq v\in\mathbb{Z}^{u_{k}}$ be the vector whose elements are the coefficients of the polynomial $(x-y)^{\ell}$, ordered following the lexicographic order; this is possible by condition (1). By the Toeplitz structure of $R_{k}$, that encodes a convolution, $R_{k}v$ contains those coefficients of $q(x,y):=(x-y)^{\ell}h_{d}(x,y)^{\ell}=(x^{d+1}-y^{d+1})^{\ell}$ that are associated with the powers of $y$ between $c_{k}$ and $c_{k}+u_{k-\ell d}-1$. Write, by Euclidean division, $c_{k}=q_{k}(d+1)+g_{k}$. By condition (2),
$q_{k}(d+1)<c_{k}\leq c_{k}+u_{k-\ell d}-1<q_{k}(d+1)+d+1\leq(1+q_{k})(d+1),$
ensuring $R_{k}v=0$ because the only nonzero coefficients of $q(x,y)$ are associated with the powers of $y$ that are exact multiples of $d+1$.
Example 4.17
We illustrate the concrete application of Proposition 4.16 for $d=3$ and $\ell=2$. Then, the $\gamma_{i}$ are defined as the coefficients of
$(1+z+z^{2}+z^{3})^{2}=1+2z+3z^{2}+4z^{3}+3z^{4}+2z^{5}+z^{6}.$
Take now $m=4,n=8,k=9$. Then, $\ell d+1=7\leq k\leq 11=m+n-1$ and $c_{9}=1$. Moreover
$u_{9}=u_{3}=3,\qquad R_{9}=\begin{bmatrix}\gamma_{1}&\gamma_{2}&\gamma_{3}\\ \gamma_{0}&\gamma_{1}&\gamma_{2}\\ \gamma_{-1}&\gamma_{0}&\gamma_{1}\end{bmatrix}=\begin{bmatrix}2&3&4\\ 1&2&3\\ 0&1&2\end{bmatrix}\Rightarrow\rho(9,4,8,3,2)=2<3.$
Note that the null vector is $\begin{bmatrix}1\\ -2\\ 1\end{bmatrix}$, which encodes the coefficients of $(x-y)^{2}$, and $(x-y)^{2}\in\mathcal{I}_{2}$ for the chosen values of $m,n,d,\ell$.
The idea of Proposition 4.16 can be generalized by taking, more generally, polynomials in the ideal $\langle(x-y)^{\ell}\rangle$. Example 4.18 shows one such construction at work.
Example 4.18
We illustrate one instance of a rank drop that is not predicted by Proposition 4.16. Let $\ell=1$ so that $\gamma_{i}\in\{0,1\}$. Take also $d=2,m=n=6,$ and $k=7$. Then, $u_{7}=u_{5}=5$, $c_{7}=1$ and
$R_{7}=\begin{bmatrix}1&1&0&0&0\\ 1&1&1&0&0\\ 0&1&1&1&0\\ 0&0&1&1&1\\ 0&0&0&1&1\end{bmatrix}\Rightarrow\begin{bmatrix}1\\ -1\\ 0\\ 1\\ -1\end{bmatrix}\in\ker R_{7}.$
From the viewpoint of ring theory, the vector in the null space encodes the polynomial $(x^{3}+y^{3})(x-y)$, which belongs to $\mathcal{I}_{1}$ for the chosen values of the variables, since $h_{2}(x,y)(x-y)(x^{3}+y^{3})=x^{6}-y^{6}$.
While these examples and partial results suggest that our insight of constructing polynomials in the annihilator ideals of $h_{d}(x,y)]^{\ell}$ is a powerful idea, a complete characterization of the ranks of $R_{k}$ involves the solution of the following combinatorial-algebraic problem: Given $d,\ell,m,n$ construct all the possible homogeneous polynomials $v(x,y)$ such that $v(x,y)h_{d}(x,y)^{\ell}$ has sufficiently many consecutive zero coefficients (where the meaning of “sufficiently many” is that arising from the previous discussion). We find that a full analysis is beyond the scope of the current manuscript, and hence we conclude with the open Problem 4.19, whose solution would allow us to refine the statement of Theorem 4.12.
Problem 4.19
Fully characterize the quadruples $(m,n,d,\ell)$ for which some $R_{k}$ in Definition 4.10 is rank deficient, and quantify exactly the amount of these rank drops.
Problem 4.19 is related to certain deep modern results in commutative algebra; namely, it can be seen as a generalization (from linear to non-linear polynomials in the graded Artin ring $\mathcal{R}=\mathbb{F}[x,y]/\langle x^{m},y^{n}\rangle$) of the problem of determining the membership, or lack thereof, of $h_{d}(x,y)^{\ell}]\in\mathcal{R}$ to the non-Lefschetz locus of $\mathcal{R}$ [3, Definition 2.3].
@@ 5 Conclusions and future outlook
We studied two nontrivial problems of combinatorial nature, the Fréchet-Jordan problem and the more general bivariate Jordan problem. We solved the Fréchet-Jordan problem by reducing it to computing either the Euclidean division of two numbers or to computing ranks of certain integer Toeplitz matrices. For the bivariate Jordan problem, we gave some partial results, complementing the existing literature. A full solution of the bivariate Jordan problem remains an open problem, and indeed not an easy one in the author’s opinion. Fully characterizing the ranks of the integer Toeplitz matrices that appear in our solution of the Fréchet-Jordan problem, or extending our analysis to fields of prime characteristic, are also possible directions for future research.
@@ 6 Acknowledgements
The late Nick Higham suggested the Fréchet-Jordan problem to me when I was a postdoc in his group. Nick himself, Bruno Iannazzo, and Samuel Relton (as well as Giovanni Barbarino, a few years later) shared comments on an earlier version of this document, that did not go much beyond Theorem 3.1. This earlier version was in fact submitted, and an anonymous reviewer pointed me to [10], which in turn cites [12]. I retracted that version, and decided to resubmit after some time and after having obtained some more results.